# Environmentally Sustainable Software Design and Development: A Systematic Literature Review

### v1

@@ 4.1. Who?
From Figure 2, there is an evident unbalance between design- and development-related stakeholders. Indeed, software developers are by far the most frequent stakeholders, being them targeted by the vast majority of selected studies (i.e., 45 out of the 65 selected studies). The frequency of software developers is also significantly higher than those of software design-related stakeholders (i.e., software architects and sustainability engineers), even if they are summed together. Given that our focus is on the design and development of sustainable software, we would have expected a more balanced distribution among design- and development-related stakeholders. In this perspective, and to avoid putting all the burden of environmental sustainability on software developers, there is a need for more solutions targeting software architects and sustainability engineers, who should be better supported in the design of environmentally sustainable software.
Other interesting observations come from the coverage of stakeholders that are not directly participating in software design and development, like end users, product owners, IT operators, and software providers. While the frequency of IT operators may be justified by the emergent wave of DevOps practices (Azad and Hyrynsalmi, 2023), the frequencies of the other stakeholders seem to suggest that the design and development of environmentally sustainable software naturally involves more stakeholders than just those involved in the design and development of software. This holds especially for end users and product owners, whose frequencies are almost the same as those of design-related stakeholders, since they can concretely impact on the design and development of environmentally sustainable software. Indeed, environmental sustainability can be a discriminant for the software chosen by end users (Jayanthi et al., 2021; Lago and Jansen, 2010; Atkinson et al., 2014; Mancebo et al., 2021). Additionally, the choices taken by product owners along the software process can affect the environmental sustainability of the software itself (Lami et al., 2014; Moshnyaga, 2013a; Sikand et al., 2023; Moshnyaga, 2013c). The above are just first insights on how different stakeholders impact on the design and development of environmentally sustainable software, even if not directly participating in software design and development themselves. Such insights shed light on the need for further investigation on the roles, contributions, and support needed by all the possible stakeholders for the design and development of environmentally sustainable software.
@@
The design and development of environmentally sustainable software involves multiple different stakeholders, whose roles, contributions, and needed support should be further investigated. |
@@ 4.2. What?
Figure 3 displays the frequency of proposals in the selected studies, clearly outlining a predominance of reference models, proposed in 23 out of 65 of the selected studies. Out of these, six studies are centered around GREENSOFT, a reference model first proposed by Naumann et al. (Naumann et al., 2011), which was then retaken and extended in five subsequent studies (Dick et al., 2013; Johann et al., 2011; Dick et al., 2010; Mahmoud and Ahmad, 2013; Naumann et al., 2015). The remaining efforts propose different reference models, however still sharing some commonalities with GREENSOFT and with each other. Whilst the number of reference models witnesses the need for clarifying what is environmental sustainability for software design and development, the lack of a commonly accepted reference model calls for standardisation efforts along this line.
@@
The existing proposals for designing and developing environmentally sustainable software then mainly focus on providing a first support to the purpose, namely sharing guidelines in the form of good practices or providing estimate techniques to assess the environmental sustainability of the software under design/development. The rest is then scattered among different attempts (especially, software improvement techniques) aimed at supporting the design and development of environmentally sustainable software, however without any concrete proposal emerging as the most promising or used for the task. Further research is, therefore, needed to establish which techniques can effectively support designing and developing environmentally sustainable software.
@@
On another front, Figure 4a shows that the vast majority of the existing proposals adopt a black box approach. Namely, most of the existing proposals do not need the source code of a software to be available. This is mainly motivated by the above distribution of proposals, since reference models and good practices (which are the most frequent categories) typically abstract from the source code of a software. Another motivation comes from the black-box software improvement techniques (Karunakaran and Rao, 2013; Heithoff et al., 2023; Beghoura et al., 2014; Jebraeil et al., 2017; Alofi et al., 2022, 2023; Tee et al., 2014a; Wei et al., 2023; Zhang et al., 2023; Lago, 2019; Lasla et al., 2022; Koçak and Alptekin, 2019), which choose a black-box approach to increase their applicability, namely to work also with software whose sources are not available, e.g., for privacy or legacy reasons. At the same time, the existing white-box proposals show concrete potentials to enhance the environmental sustainability of the software under design/development, as witnessed, e.g., by the green code smells detected by the ecoCode project (Goaër, 2021, 2023; Goaër and Hertout, 2022). Following this line, it may be worthy to consider further investigating the potentials of white-box analysis, as well as finding trade-offs between proposals’ applicability and enhancements in terms of sustainability.
@@
Finally, Figure 4b displays the overall low TRL of the state of the art on designing and developing environmentally sustainable software. Most of the studies feature a TRL of 2 or 3, which – together with TRL 1 – cover around 66% of the selected studies. Among those, only 21.5% of the selected studies went beyond observing the basic principles and formulating the technology concepts, by proposing an experimental proof-of-concept. 32.3% of the selected studies went a step higher, by validating their proposal in laboratory and, therefore, reaching TRL 4. Only the experience report by Mehra et al. (Mehra et al., 2023) reached TRL 5, being it a report on a industry-relevant case study. Overall, the above described distribution of TRLs witnesses a need for increasing the TRL of what is available to support the design and development of environmentally sustainable software, especially if we wish to support the software industry in the task.
@@
Most of the proposals are reference models, witnessing the neeed for clarifying what is the environmentally sustainable software design and development. | |
@@
The overall TRL is low and should be increased to favour the adoption of existing solutions by the software industry. |
@@ 4.3. Why?
The why question aims at identifying the sustainability goals in the selected studies, namely whether they support the design and development of environmentally sustainable software by targeting its energy consumption or carbon footprint. As shown in Figure 5, the most frequent trends are reducing the energy consumption or carbon footprint of software.
@@
There are also solutions proposed to enable measuring energy consumption or carbon emissions, which are among the goals of ten and seven of the selected studies, respectively. The underlying idea there is in line with that of the studies targeting the reduction of energy consumption or carbon emissions. More precisely, the solutions aimed at measuring energy consumption or carbon emissions are intended to enable comparing the energy consumption or carbon emissions before/after updating a software, with the ultimate goal of assessing whether any applied or planned change contributes to reducing them (Patterson et al., 2022; Kern et al., 2013; Lasla et al., 2022; Kipp et al., 2011).
@@
Only three selected studies (Alofi et al., 2023, 2022; Zhang et al., 2023) pursue the goal of minimising both energy consumption or carbon footprint of software. Such a low number is not only motivated by the type of proposals, as highlighted above, but also by the inherent complexity of optimising by considering the multiple different aspects that influence the energy consumption and carbon footprint of the software under design/development (Mancebo et al., 2021; Siegmund et al., 2022; Wedyan et al., 2023; Tee et al., 2014a; Mehra et al., 2022). Indeed, Aloifi et al. (Alofi et al., 2022, 2023) and Zhang et al. (Zhang et al., 2023) pursued their optimisation goals by narrowing the search space by relying on the context given by blockchains and ML, respectively. These are quite recent efforts, which anyhow showed the feasibility of minimising a software’s energy consumption and carbon footprint. Further investigation is needed along this line, possibly also considering, e.g., heuristics to achieve sub-optimal minimisation, if the holistic minimisation would get too complex to achieve.
@@
Other interesting observations come from the Figure 6, which shows that only seven out of the 65 selected studies consider only the carbon footprint of software. Indeed, most of the selected studies consider the energy consumption of software as the main discriminant to pursue their environmental sustainability goals. This is done by only considering the energy consumption of software (in 18 selected studies), or by considering it in conjunction with its carbon footprint (in 40 selected studies). In the latter case, the carbon footprint is typically estimated based on the energy consumption, to then consider both while pursuing the goal of measuring/reducing/minimising both energy consumption and carbon footprint. This is mainly motivated by an assumption shared by the selected studies considering energy consumption, namely that reducing such energy consumption of software also reduces its carbon footprint. However, as demonstrated by Aiello et al. (Aiello et al., 2024), energy mixes play a crucial role to relate energy consumption with carbon footprint, as different energy mixes come with different associated carbon intensities. This may negatively impact on the effectiveness of directly considering energy consumption to indirectly reduce carbon emission. For example, when a software is powered with clean and renewable energy, there may be no need for reducing its energy consumption, as this may not significantly change its carbon footprint (Forti et al., 2024). This calls for further research along the line of directly considering energy mixes and carbon footprint, rather than as an indirect consequence of energy consumption.
@@
The most pursued goals are reducing energy consumption and/or carbon emissions. Their minimisation can however be achieved by relying on the application domain to narrow the search space of the minimisation problem. | |
Using energy consumption as the main discriminant may not be enough, as different energy mixes may result in different carbon emissions. This calls for solutions directly considering energy mixes and carbon footprint as well. |
@@ 4.4. Where?
Figure 7 displays the frequency of application domains in the selected studies, clearly outlining that the vast majority of existing solutions target generic software – rather than application domain-specific software.
This is mainly motivated by the fact that sustainability guidelines, reference models, and measurement solutions are given as high-level solutions to reduce energy consumption and/or carbon emissions of generic software (Figure 8), possibly requiring some adaptation for getting applied to application domain-specific software. The applicability and extension of generic reference models to specific application domains is, therefore, deserving further investigation.
@@
Narrowing the scope to specific application domains enables providing more concrete techniques to design and develop environmentally sustainable software, calling for further investigation along this line. |
@@ 4.5. When?
Figure 9 shows the frequency of targeted lifecycle stages in the selected studies.
Being the present SLR targeting the design and development of environmentally sustainable software, such stages naturally result to be the most targeted by the selected studies. The other three stages are anyhow quite frequent in the selected studies, with requirement analysis targeted by 40 studies, integration & testing targeted by 34 studies, and rollout targeted by 33 studies. This is mainly motivated by the fact that the currently available reference models and measurement solutions, which together constitute more than a half of the selected studies, span over all lifecycle stages (Figure 10a).
Differently, sustainability guidelines and software improvement techniques seem to be more focused on the design and develepment stages, with less than a half of them targeting other lifecycle stages (Figure 10).
On another front, it is interesting to observe who is involved in which lifecycle stage, according to the state of the art on the design and development of environmentally sustainable software. Figure 10b provides such information, confirming what we observed in Section 4.1: the most involved stakeholders are software developers. Notably, software developers are not only involved in the development stage, but rather result the most involved across all the stages of a software’s lifecycle – which is again in line with what we observed in in Section 4.1, namely that the burden of environmental sustainability is expected to be put on software developers. This, along with the fact that the spanning across all stages seem to apply to all stakeholders, suggests that all stakeholders should be involved – and supported – during all lifecycle stages. Therefore, further investigation should be pursued along this line.
@@
Software developers are the most involved in all stages, but all stakeholders span across all lifecycle stages, calling for supporting all stakeholders during all lifecycle stages. |
@@ 4.6. How?
Notably, the peak in 2013 and the ongoing growth seem to relate two co-occurring events in time. More precisely, the peak in 2013 may be a consequence of a report (OECD, 2009) by OECD (Organisation for Economic Cooperation and Development), which assessed government programmes and business initiatives on ICT and environmental sustainability across 22 OECD countries. Instead, the ongoing growth since 2020 may be a consequence of the recent interests in environmental sustainability. These were raised by, e.g., by international organisations like Fridays for Future, and investment policies like the European strategic agenda (European Union, 2019b), which calls and finances a sustainable growth of European industries, therein included the IT industry (European Union, 2019a, 2021, 2024).
@@
An ever increasing increasing number of publications since 2020 witnesses the hotness of the topic, also pairing with the ever increasing interest in environmental sustainability. |
@@ 5. Threats to Validity
Threats to External Validity. The potential threats to the external validity of a study affect the applicability of the study’s results in a broader and more general context (Wohlin et al., 2000). In our case, since the selected studies are obtained from a large extent of online available sources, our results and observations may be only partly applicable to the broad area of the design and development of environmentally sustainable software, and this may potentially threaten the external validity of our study. To mitigate this potential threat, we organised multiple feedback sessions during our study. We analysed the discussion following up from each feedback session, and we exploited the resulting qualitative data to fine-tune both our research methods and the applicability of our findings. We also prepared a replication package bundling the intermediary artefacts and the final results produced while running our study, which is publicly available online (Danushi et al., 2024). The replication package, together with the references to the selected studies listed in Table 2, can be used to deepen the understanding on our data, as well as to repeat of our study and verify of our findings.
@@
Threats to Internal Validity The internal validity of a study concerns the validity of the method employed to analyse the study data (Wohlin et al., 2000). To reinforce the internal validity of our study, we relied on established guidelines to conduct systematic literature reviews in software engineering (Kitchenham and Charters, 2007; Petersen et al., 2008; Petersen et al., 2015). Additionally, we enacted various triangulation rounds and adopted inter-rater reliability assessment (see Section 2), which were conceived to avoid possible biases by triangulation.
Threats to Construct and Conclusions Validity The construct and conclusions validities concerns the generalisability of the constructs under study and the degree to which the conclusions of a study are reasonably based on the available data, respectively (Wohlin et al., 2000). The construct and conclusions validities of our study may have been affected by observer and interpretation biases. To mitigate these potential threats, we exploited inter-rater reliability assessment (Section 2), and we also independently drawn conclusions from the available data. Such conclusions were then suitably merged and double-checked against the selected (and related) studies in a joint discussion session.

### v2

@@ 4.1. Who?
From Figure 2, there is an evident unbalance between design- and development-related stakeholders. Indeed, software developers are by far the most frequent stakeholders, who are targeted by the vast majority of selected studies (i.e., 45 out of the 65 selected studies). The frequency of software developers is also significantly higher than that of software design-related stakeholders (i.e., software architects and sustainability engineers), even when added together. Given that our focus is on the design and development of sustainable software, we would have expected a more balanced distribution among design- and development-related stakeholders. In this perspective, and to avoid putting all the burden of environmental sustainability on software developers, there is a need for more solutions targeting software architects and sustainability engineers, who should be better supported in the design of environmentally sustainable software.
Other interesting observations come from the coverage of stakeholders not directly involved in software design and development, like end users, product owners, IT operators, and software providers. While the frequency of IT operators may be justified by the emergent wave of DevOps practices (Azad and Hyrynsalmi, 2023), the frequencies of the other stakeholders seem to suggest that the design and development of environmentally sustainable software naturally involves more stakeholders than just those involved in the design and development of software. This holds especially for end users and product owners, whose frequencies are almost the same as those of design-related stakeholders, since they can concretely impact on the design and development of environmentally sustainable software. Indeed, environmental sustainability can be a discriminant for the software chosen by end users (Jayanthi et al., 2021; Lago and Jansen, 2010; Atkinson et al., 2014; Mancebo et al., 2021). Additionally, the choices taken by product owners along the software process can affect the environmental sustainability of the software itself (Lami et al., 2014; Moshnyaga, 2013a; Sikand et al., 2023; Moshnyaga, 2013c). The above are just first insights on how different stakeholders impact on the design and development of environmentally sustainable software, even if not directly participating in software design and development themselves. Such insights shed light on the need for further investigation on the roles, contributions, and support needed by all the possible stakeholders for the design and development of environmentally sustainable software.
@@
| The design and development of environmentally sustainable software involves multiple different stakeholders, whose roles, contributions, and needed support should be further investigated. | |
@@ 4.2. What?
Figure 3 displays the frequency of proposals in the selected studies, clearly outlining a predominance of reference models, proposed in 23 out of the 65 selected studies. Of these, six studies focus on GREENSOFT, a reference model first proposed by Naumann et al. (Naumann et al., 2011) and then rivisited and extended in five subsequent studies (Dick et al., 2013; Johann et al., 2011; Dick et al., 2010; Mahmoud and Ahmad, 2013; Naumann et al., 2015). The remaining efforts propose different reference models, however still sharing some commonalities with GREENSOFT and with each other. Whilst the number of reference models witnesses the need for clarifying what is environmental sustainability for software design and development, the lack of a commonly accepted reference model calls for standardisation efforts along this line.
@@
The existing proposals for designing and developing environmentally sustainable software then mainly focus on providing a first support to the purpose, namely sharing guidelines in the form of good practices or providing estimate techniques to assess the environmental sustainability of the software under design/development. The rest is then scattered among different attempts (especially, software improvement techniques) to support the design and development of environmentally sustainable software, however without any concrete proposal emerging as the most promising or used for the task. Further research is, therefore, needed to establish which techniques can effectively support designing and developing environmentally sustainable software.
@@
On another front, Figure 4a shows that the vast majority of the existing proposals adopt a black box approach. Namely, most of the existing proposals do not need the source code of software to be available. This is mainly motivated by the above distribution of proposals, since reference models and good practices (which are the most frequent categories) typically abstract from the source code of a software. Another motivation comes from the black-box software improvement techniques (Karunakaran and Rao, 2013; Heithoff et al., 2023; Beghoura et al., 2014; Jebraeil et al., 2017; Alofi et al., 2022, 2023; Tee et al., 2014a; Wei et al., 2023; Zhang et al., 2023; Lago, 2019; Lasla et al., 2022; Koçak and Alptekin, 2019), which choose a black-box approach to increase their applicability, namely to work also with software whose sources are not available, e.g., for privacy or legacy reasons. At the same time, the existing white-box proposals show concrete potentials to enhance the environmental sustainability of the software under design/development, as witnessed, e.g., by the green code smells detected by the ecoCode project (Goaër, 2021, 2023; Goaër and Hertout, 2022). Following this line, it may be worthy to consider further investigating the potentials of white-box analysis, as well as finding trade-offs between proposals’ applicability and enhancements in terms of sustainability.
@@
Finally, Figure 4b displays the overall low TRL of the state of the art on designing and developing environmentally sustainable software. Most of the studies feature a TRL of 2 or 3, which – together with TRL 1 – cover around 66% of the selected studies. Among those, only 21.5% of the selected studies went beyond observing the basic principles and formulating the technology concepts, by proposing an experimental proof-of-concept. 32.3% of the selected studies went a step higher, by validating their proposal in laboratory and, therefore, reaching TRL 4. Only the experience report by Mehra et al. (Mehra et al., 2023) reached TRL 5, being it a report on an industry-relevant case study. Overall, the above described distribution of TRLs witnesses a need for increasing the TRL of what is available to support the design and development of environmentally sustainable software, especially if we wish to support the software industry in this task.
@@
| Most of the proposals are reference models, witnessing the need for clarifying what is the environmentally sustainable software design and development. | |
@@
| The overall TRL is low and should be increased to favour the adoption of existing solutions by the software industry. | |
@@ 4.3. Why?
The why question aims at identifying the sustainability goals in the selected studies, namely whether they support the design and development of carbon-efficient software by directly or indirectly considering its carbon footprint. As shown in Figure 5, the most frequent trends are reducing the carbon footprint of software through a hybrid approach, namely directly considering a software’s carbon footprint, as well as considering the indirect effect of its energy consumption on such footprint.
@@
There are also solutions proposed to enable directly/indirectly measuring carbon emissions, which are among the goals of ten and seven of the selected studies, respectively. The underlying idea is in line with that of the studies targeting the reduction of carbon emissions. More precisely, the solutions aimed at directly/indirectly measuring carbon emissions are intended to enable comparing the carbon emissions before/after updating a software, with the ultimate goal of assessing whether any applied or planned change contributes to reducing them (Patterson et al., 2022; Kern et al., 2013; Lasla et al., 2022; Kipp et al., 2011).
@@
Only three selected studies (Alofi et al., 2023, 2022; Zhang et al., 2023) pursue the goal of minimising the carbon footprint of software, doing this with a hybrid approach. Such a low number is not only motivated by the type of proposals, as highlighted above, but also by the inherent complexity of optimising by considering the multiple different aspects that influence the carbon footprint of the software under design/development (Mancebo et al., 2021; Siegmund et al., 2022; Wedyan et al., 2023; Tee et al., 2014a; Mehra et al., 2022). Indeed, Alofi et al. (Alofi et al., 2022, 2023) and Zhang et al. (Zhang et al., 2023) pursued their optimisation goals by narrowing the search space by relying on the context given by blockchains and ML, respectively. These are quite recent efforts, which anyhow showed the feasibility of minimising a software’s carbon footprint. Further investigation is needed along this line, possibly also considering, e.g., heuristics to achieve sub-optimal minimisation, if the holistic minimisation would get too complex to achieve.
@@
Other interesting observations come from Figure 6, which shows that only seven out of the 65 selected studies consider only the carbon footprint of software. Indeed, most of the selected studies pursue their carbon efficiency goals with an indirect or hybrid approach (in 18 and 40 of the selected studies, respectively). In both cases, the carbon footprint is estimated starting from energy consumption while pursuing the goal of measuring/reducing/minimising a software’s carbon footprint. This is mainly motivated by an assumption shared by the selected studies considering energy consumption, namely that reducing the energy consumption of software also reduces its carbon footprint. However, as demonstrated by Aiello et al. (Aiello et al., 2024), energy mixes play a crucial role to relate energy consumption with carbon footprint, as different energy mixes have different associated carbon intensities. This may negatively impact on the effectiveness of considering energy consumption to indirectly reduce carbon emission. For example, when a software is powered with clean and renewable energy, there may be no need for reducing its energy consumption, as this may not significantly change its carbon footprint (Forti et al., 2024). This calls for further research along the line of directly considering energy mixes and carbon footprint, rather than as an indirect consequence of energy consumption.
@@
| The most pursued goals are reducing carbon emissions. Their minimisation can anyhow be achieved by relying on the application domain to narrow the search space of the minimisation problem. | |
| Reducing energy consumption to indirectly reduce a software’s carbon footprint may not be enough, as different energy mixes may result in different carbon emissions. This calls for solutions directly considering energy mixes and carbon footprint as well. | |
@@ 4.4. Where?
Figure 7 displays the frequency of application domains in the selected studies, outlining that the vast majority of existing solutions target generic software – rather than application domain-specific software.
This is mainly motivated by the fact that sustainability guidelines, reference models, and measurement solutions are given as high-level, broadly-applicable approaches to reduce the carbon emissions of software, as it can be observed from the heatmaps in Figure 8. The applicability and extension of such generic approaches to specific application domains is, therefore, deserving further investigation, as they might require some adaptation for getting applied to application domain-specific software. The same does not hold for the available software improvement techniques, as most of them target specific application domains (Figure 8), rather exploiting the peculiarities of application domain-specific software for reducing its carbon emissions.
@@
| Narrowing the scope to specific application domains enables providing more concrete techniques to design and develop carbon-efficient software, calling for further investigation along this line. | |
@@ 4.5. When?
Figure 9 shows the frequency of lifecycle stages targeted in the selected studies.
Being the present SLR targeting the design and development of environmentally sustainable software, such stages naturally result to be the most targeted by the selected studies. The other three stages are anyhow quite frequent in the selected studies, with requirement analysis targeted by 40 studies, integration & testing targeted by 34 studies, and rollout targeted by 33 studies. This is mainly due to the fact that the currently available reference models and measurement solutions – which together account for more than half of the selected studies – are designed to be applicable across all lifecycle stages, thereby naturally extending their relevance beyond design and development.
In contrast, sustainability guidelines and software improvement techniques tend to be more concentrated on these two stages, with less than half addressing other phases of the lifecycle (Figure 10).
On another front, it is interesting to observe who is involved in which lifecycle stage, according to the state of the art on the design and development of environmentally sustainable software. Figure 10b provides such information, confirming what we observed in Section 4.1: the most involved stakeholders are software developers. Notably, software developers are not only involved in the development stage, but rather result the most involved across all the stages of a software’s lifecycle – which is again in line with what we observed in in Section 4.1, namely that the burden of environmental sustainability is expected to be put on software developers. This, along with the fact that the spanning across all stages seems to apply to all stakeholders, suggests that all stakeholders should be involved – and supported – during all lifecycle stages. Therefore, further investigation should be pursued along this line.
@@
| Software developers are the most involved in all stages, but all stakeholders span across all lifecycle stages, calling for supporting all stakeholders during all lifecycle stages. | |
@@ 4.6. How?
Notably, the peak in 2013 and the ongoing growth seem to relate two co-occurring events in time. More precisely, the peak in 2013 may be a consequence of a report (OECD, 2009) by OECD (Organisation for Economic Cooperation and Development), which assessed government programmes and business initiatives on ICT and environmental sustainability across 22 OECD countries. Instead, the ongoing growth since 2020 may be a consequence of the recent interest in environmental sustainability. These were raised, e.g., by international organisations like Fridays for Future, and investment policies like the European strategic agenda (European Union, 2019b), which calls for and finances a sustainable growth in European industries, including the IT industry (European Union, 2019a, 2021, 2024).
@@
| An ever increasing number of publications since 2020 witnesses the hotness of the topic, also pairing with the ever increasing interest in environmental sustainability. | |
@@ 5. Threats to Validity
Threats to External Validity. The potential threats to the external validity of a study affect the applicability of the study’s results in a broader and more general context (Wohlin et al., 2000). In our case, since the selected studies are obtained from a large extent of online available sources, our results and observations may be only partly applicable to the broad area of the design and development of carbon efficient software, and this may potentially threaten the external validity of our study. To mitigate this potential threat, we organised multiple feedback sessions during our study. We analysed the discussion following up from each feedback session, and we exploited the resulting qualitative data to fine-tune both our research methods and the applicability of our findings. We also prepared a replication package bundling the intermediary artefacts and the final results produced while running our study, which is publicly available online (Danushi et al., 2025). The replication package, together with the references to the selected studies listed in Table 2, can be used to deepen the understanding of our data, as well as to repeat our study and verify our findings.
@@
Threats to Internal Validity The internal validity of a study concerns the validity of the method employed to analyse the study data (Wohlin et al., 2000). To reinforce the internal validity of our study, we relied on established guidelines to conduct systematic literature reviews in software engineering (Kitchenham and Charters, 2007; Petersen et al., 2008, 2015). Additionally, we enacted various triangulation rounds and adopted inter-rater reliability assessment (see Section 2), which were conceived to avoid possible biases by triangulation.
Threats to Construct and Conclusions Validity The construct and conclusions validities concerns the generalisability of the constructs under study and the degree to which the conclusions of a study are reasonably based on the available data, respectively (Wohlin et al., 2000). The construct and conclusions validities of our study may have been affected by observer and interpretation biases. To mitigate these potential threats, we exploited inter-rater reliability assessment (see Section 2), and we also independently drawn conclusions from the available data. Such conclusions were then suitably merged and double-checked against the selected (and related) studies in a joint discussion session.